OVHCloud Web Hosting Status

Current status
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#10110 — LNS-1-PAR
Incident Report for Web Cloud
SFR collect on lns-1-par is down, we are investigating it with the manufacturer.
The customers were automatically switched onto the backup LNS.


Date: 2014-02-06 14:59:12 UTC
We have received the necessary patch and we will update the LNS tonight at around 0:00

Date: 2014-01-31 16:58:22 UTC
We found a bug on the Ericsson router, we will receive the patch next week.
We will migrate the remaining FT customers onto the Roubaix LNS, and we will upgrade the node when we get the patch.

Date: 2014-01-31 10:40:58 UTC
We carried out tests with SFR without managing to find the anomaly, we are now working with Ericsson to find the source of the problem.

Date: 2014-01-30 10:44:59 UTC
We have changed the main LNS configuration LNS for lns-1-par customers to lns-1-lyo with lns-1-rbx as a backup, so that the sessions don't come back onto the Paris LNS, which is causing the instabilities.
The PPP sessions will thus be the Lyon LNS at the time of the next reset.

Date: 2014-01-30 09:34:55 UTC
Following the return of the SFR collect to the Paris LNS during the night, instabilities have been detected.
The port was shutdown, access is on the backup LNS again.
Posted Jan 29, 2014 - 13:21 UTC